A stream of water vapor and air has a dew point of 60C. If the total pressure is 760 mm
of Hg, calculate the mole fraction of oxygen in the stream.
1
Expert's answer
2020-07-07T14:39:04-0400

the actual vapor pressure (p) can be

calculated using the formula listed below:

p= "6.11*10^{(7.5*T_d)\/(237.3+T_d)}"

where Td is dew point and unit of p is millibar

so p at dew point is

p= "6.11*10^{7.5*60\/(237.3+60)}"

so p = 199.3 milli bar

and 1 millibar = 0.75 mm Hg

so p will be 149.5 mm Hg


so mole fraction will be p/Po

where Po = 760 mm Hg

so mole fraction = 149.5/760 = 0.196

this is the mole fraction of water vapour.


mole fraction of oxygen = 1-0.196 = 0.804